OpenStudy (anonymous):

∫ln(2) to 0 e^x sin(1/2e^(2x) π)dx calculates out to: FresnelS(2) - FresnelS( 1) remebering that that FresnelS(x)=∫0 to x sin((π t^2)/2)dt use a transformation to explain this calcluation. explain in detail how you go through it. please

OpenStudy (experimentx):

could you use latex? or draw app http://www.codecogs.com/latex/eqneditor.php

OpenStudy (experimentx):

do you mean \[ \int_{\ln 2}^{0} e^{x} \sin \left( \frac{1}{2}e^{2x} \pi \right) dx\]

OpenStudy (anonymous):

My literacy sheet show it the other way, 0 on bottom